# Optimal Frames for Phase Retrieval from Edge Vectors of Optimal Polygons

### v2

@@ 1.3.3. Characterization of all optimal frames
Theorem 1.4 establishes that any optimal frame ${\mathcal{A}}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset{\mathbb{R}}^{2}$ forms a tight frame when $m\geq 3$ has an odd factor. Inspired by this result, we propose the following conjecture: Any optimal frame ${\mathcal{A}}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset{\mathbb{R}}^{d}$ for phase retrieval in ${\mathbb{R}}^{d}$ must be a tight frame, provided $m\geq 2d-1$.
Remark 1.6.
Let $m\geq 3$ be an integer with an odd factor. A brute-force algorithm can be used to generate all optimal polygons in ${\mathcal{E}}_{m}$, up to scaling, translation, rotation, and flip. It can also generate all optimal frames in $\mathcal{T}_{m}$, up to scaling, rotation, and flip  [^1] The MATLAB and Maple implementations of the algorithm are available at https://github.com/zxynhy/Optimal-polygon-for-the-perimeter-maximizing-isodiametric-problem.git.. The algorithm begins by examining all vectors $\boldsymbol{\varepsilon}\in\{\pm 1\}^{m}$ that satisfy (1.11). Then Theorem 1.3 and Theorem 1.4 are applied. Take $m=12$ as an example. After checking all vectors $\boldsymbol{\varepsilon}\in\{\pm 1\}^{12}$ for which (1.11) holds, by Theorem 1.3, we identify all optimal polygons in ${\mathcal{E}}_{12}$ (up to scaling, translation, rotation and flip), shown in Figure 2 (a1) and (a2). Also, by Theorem 1.4, all the optimal frames in $\mathcal{T}_{12}$ (up to scaling, rotation and flip) are obtained, shown in Figure 2 (b1) and (b2). Denote by $\mathcal{T}_{m}$ the number of distinct frames in $\mathcal{T}_{m}$, modulo scaling, rotation and flip. Table 1 lists the number $\mathcal{T}_{m}$ for $3\leq m\leq 15$. In particular, the cases $\mathcal{T}_{4}=\#\mathcal{T}_{8}=1$ follow from established results in [29, 28, 4, 9]. A promising direction for future research is to design an efficient algorithm capable of finding all solutions to Problem 1.2.
Table 1. The number $\mathcal{T}_{m}$ of distinct frames in $\mathcal{T}_{m}$, modulo scaling, rotation and flip, for $m\in[3,15]$.
$m$ 3 4 5 6 7 8 9 10 11 12 13 14 15
$\mathcal{T}_{m}$ 1 1 1 1 1 1 2 1 1 2 1 1 5
@@ 2.1. Notation
For a positive integer $m$, we define $m]:=\{1,\dotsc,m\}$. Let $\mathbb{H}\in\{\mathbb{R},\mathbb{C}\}$. We use $\|\boldsymbol{x}\|_{2}$ to denote the Euclidean norm of a vector $\boldsymbol{x}\in\mathbb{H}^{d}$, and we use $\|\boldsymbol{A}\|_{2}$ to denote the spectral norm of a matrix $\boldsymbol{A}\in\mathbb{H}^{m\times d}$. We denote the unit sphere in $\mathbb{H}^{d}$ by $\mathbb{S}_{\mathbb{H}}^{d-1}$, i.e.,
@@
2.2. Equivalent forms of Lipschitz constants and condition numbers
@@
The following lemma yields an equivalent form for the condition number of a tight frame.
@@
$\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})$ $=\min_{c\in\mathbb{H},|c|=1}\|\boldsymbol{x}-c\cdot\boldsymbol{y}\|_{2}$
@@
$L_{\boldsymbol{B}}$ $=\inf_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})\neq 0\end{subarray}}\frac{\||\boldsymbol{B}\boldsymbol{x}|-|\boldsymbol{B}\boldsymbol{y}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})}$
$\geq\frac{1}{\|\boldsymbol{M}^{\frac{1}{2}}\|_{2}}\cdot\inf_{\begin{subarray}{c}\widehat{\boldsymbol{x}},\widehat{\boldsymbol{y}}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\widehat{\boldsymbol{x}},\widehat{\boldsymbol{y}})\neq 0\end{subarray}}\frac{\||\boldsymbol{A}\widehat{\boldsymbol{x}}|-|\boldsymbol{A}\widehat{\boldsymbol{y}}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\widehat{\boldsymbol{x}},\widehat{\boldsymbol{y}})}=\frac{L_{\boldsymbol{A}}}{\|\boldsymbol{M}^{\frac{1}{2}}\|_{2}}.$
@@
We arrive at our conclusion.
2.3. Basic properties of convex polygons
Here we introduce some basic properties of convex polygons.
@@ 2.3. Basic properties of convex polygons
[26] The width function of a convex polygon ${P}\subset\mathbb{R}^{2}$ in the direction $\boldsymbol{u}\in\mathbb{S}^{1}$ is defined by $w({P},\boldsymbol{u}):=h({P},\boldsymbol{u})+h({P},-\boldsymbol{u})$, where $h({P},\boldsymbol{z}):=\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{z}\rangle$ for any $\boldsymbol{z}\in\mathbb{S}^{1}$.
@@
$\operatorname{diam}({P})=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}w({P},\boldsymbol{u}).$
@@
$\operatorname{diam}(P)=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}w(P,\boldsymbol{u})$ $=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}(h(P,\boldsymbol{u})+h(P,-\boldsymbol{u}))$ (2.10)
$=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}(\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{u}\rangle+\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},-\boldsymbol{u}\rangle)$
$=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}(\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{u}\rangle-\min_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{u}\rangle)$
$=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}\,{\langle\boldsymbol{v}_{j_{2}}-\boldsymbol{v}_{j_{1}},\boldsymbol{u}\rangle}.$
@@
$\langle\boldsymbol{v}_{j_{2}}-\boldsymbol{v}_{j_{1}},\boldsymbol{u}\rangle=\langle\sum_{j=j_{1}}^{j_{2}-1}{\boldsymbol{e}_{j}},\boldsymbol{u}\rangle$ $\overset{(a)}{=}\langle\frac{1}{2}\sum_{j=j_{1}}^{j_{2}-1}{\boldsymbol{e}_{j}}-\frac{1}{2}\sum_{j=1}^{j_{1}-1}{\boldsymbol{e}_{j}}-\frac{1}{2}\sum_{j=j_{2}}^{m}{\boldsymbol{e}_{j}},\boldsymbol{u}\rangle$ (2.12)
$\overset{(b)}{=}\frac{1}{2}\sum_{j=1}^{m}{|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|},$
@@
$\langle\boldsymbol{e}_{j_{1}},\boldsymbol{u}\rangle=\langle\boldsymbol{v}_{j_{1}+1},\boldsymbol{u}\rangle-\langle\boldsymbol{v}_{j_{1}},\boldsymbol{u}\rangle\geq 0,\;\,\langle\boldsymbol{e}_{j_{2}-1},\boldsymbol{u}\rangle=\langle\boldsymbol{v}_{j_{2}},\boldsymbol{u}\rangle-\langle\boldsymbol{v}_{j_{2}-1},\boldsymbol{u}\rangle\geq 0.$ (2.13)
@@
$\langle\boldsymbol{e}_{t-1},\boldsymbol{u}\rangle<0,\quad\langle\boldsymbol{e}_{t},\boldsymbol{u}\rangle\geq 0,$ (2.14)
since otherwise we have $\langle\boldsymbol{e}_{i},\boldsymbol{u}\rangle<0$ for each $s+1\leq i\leq j_{2}-1$, which contradicts with (2.13). Since $P$ is strongly convex, for any $\boldsymbol{x}\in P$ there exist $\gamma_{1},\gamma_{2}\geq 0$ depending on $\boldsymbol{x}$ such that
@@
Hence, we arrive at our conclusion. This completes the proof.
@@
[21] Suppose $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}\in\mathbb{R}^{2}$ are distinct unit vectors that span $\mathbb{R}^{2}$, and suppose that $\alpha_{1},\dotsc,\alpha_{m}>0$. Then there exists a strictly convex $m$-gon P{P} in $\mathbb{R}^{2}$, having edge unit outer normals $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}$ and corresponding edge lengths $\alpha_{1},\dotsc,\alpha_{m}>0$, if and only if $\sum_{j=1}^{m}\alpha_{j}\boldsymbol{u}_{j}=\boldsymbol{0}.$ Moreover, such a polygon P{P} is unique up to translation.
@@
Assume that $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$. For all $j\in[m]$, rotating each vector $\frac{\boldsymbol{e}_{j}}{\|\boldsymbol{e}_{j}\|_{2}}$ counterclockwise by $\frac{\pi}{2}$ yields $\boldsymbol{u}_{j}$, while preserving the zero-sum property $\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}\boldsymbol{u}_{j}=\boldsymbol{0}$. It is easy to verify that $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}\in\mathbb{R}^{2}$ are distinct unit vectors that span $\mathbb{R}^{2}$. Therefore we can utilize Theorem 2.3 to obtain a strictly convex $m$-gon P{P} with edge unit outer normals $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}$ and corresponding edge lengths $\|\boldsymbol{e}_{1}\|_{2},\dotsc,\|\boldsymbol{e}_{m}\|_{2}$, i.e., with edge vectors $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$. Moreover, P{P} is unique, disregarding the translation of polygons.
Conversely, assume that P{P} is a strictly convex $m$-gon with the edge set $E$. Assume that $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$ are arranged counterclockwise. It is clear that $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$ and each $\boldsymbol{e}_{i}$ is nonzero. Without loss of generality, it is enough to prove that $\boldsymbol{e}_{1}$ and $\boldsymbol{e}_{i}$ have distinct directions for any $i>1$. For each $i\in\{2,\ldots,m\}$, define $\alpha_{i}$ as the counterclockwise angle from $\boldsymbol{e}_{1}$ to $\boldsymbol{e}_{i}$. Since $P$ is strictly convex, we have $\alpha_{i}\in(0,2\pi)$, and $\alpha_{i}$ is strictly increasing as $i$ increases. Thus, there does not exist an integer $i\in\{2,\ldots,m\}$ such that $\alpha_{i}=0$, i.e., $\boldsymbol{e}_{1}$ and $\boldsymbol{e}_{j}$ have the same direction. This completes the proof.
3. Proofs of Theorem 1.2 and Corollary 1.1
3.1. The relationship between tight frames and convex polygons in $\mathbb{R}^{2}$
@@
Let $\mathcal{A}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathbb{R}^{d}$. If there exist two distinct integers $i,j\in[m]$ such that $\boldsymbol{a}_{i}$ is a scalar multiple of $\boldsymbol{a}_{j}$, i.e., $\boldsymbol{a}_{i}=\lambda\boldsymbol{a}_{j}$ for some $\lambda\in\mathbb{R}$, then $\mathcal{A}$ is called reducible; otherwise, $\mathcal{A}$ is called irreducible.

### v3

@@ 1.2. Our contribution
Let $m\geq 3$ be an integer with an odd factor. A brute-force algorithm can be used to generate all optimal frames in $\mathcal{T}_{m}$, up to scaling, rotation, and reflection  [^1] The MATLAB and Maple implementations of the algorithm are available at https://github.com/zxynhy/Optimal-frames-for-the-stability-of-phase-retrieval.git.. The algorithm begins by determining all vectors $\boldsymbol{\varepsilon}\in\{\pm 1\}^{m}$ that satisfy (1.8). Then Theorem 1.2 is applied. Take $m=12$ as an example. After identifying all vectors $\boldsymbol{\varepsilon}\in\{\pm 1\}^{12}$ for which (1.8) holds, by Theorem 1.2, we obtain all the optimal frames in $\mathcal{T}_{12}$ (up to scaling, rotation and reflection), shown in Figure 2 (b1) and (b2). Denote by $\mathcal{T}_{m}$ the number of distinct frames in $\mathcal{T}_{m}$, modulo scaling, rotation and reflection. Table 1 lists the number $\mathcal{T}_{m}$ for integers $m\in[3,15]$ with an odd factor. A promising direction for future research is to design an efficient algorithm capable of finding all vectors $\boldsymbol{\varepsilon}\in\{\pm 1\}^{m}$ that fulfill (1.8).
Table 1. Number $\mathcal{T}_{m}$ of inequivalent optimal frames (up to scaling, rotation, and reflection) for integers $3\leq m\leq 15$ with an odd prime factor.
$m$ 3 5 6 7 9 10 11 12 13 14 15
$\mathcal{T}_{m}$ 1 1 1 1 2 1 1 2 1 1 5
Theorem 1.2 establishes that any optimal frame ${\mathcal{A}}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset{\mathbb{R}}^{2}$ forms a tight frame when $m\geq 3$ has an odd factor. Inspired by this result, we propose the following conjecture.
Conjecture 1.2.
Any optimal frame ${\mathcal{A}}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset{\mathbb{R}}^{d}$ for phase retrieval in ${\mathbb{R}}^{d}$ must be a tight frame, provided $m\geq 2d-1$.
@@ 2.1. Notation
For a positive integer $m$, we define $m]:=\{1,\dotsc,m\}$. We use $\|\boldsymbol{x}\|_{2}$ to denote the Euclidean norm of a vector $\boldsymbol{x}\in\mathbb{H}^{d}$, and we use $\|\boldsymbol{A}\|_{2}$ to denote the spectral norm of a matrix $\boldsymbol{A}\in\mathbb{H}^{m\times d}$. We denote the unit sphere in $\mathbb{H}^{d}$ by $\mathbb{S}_{\mathbb{H}}^{d-1}$, i.e.,
@@
2.2. Characterizations of Lipschitz constants and condition numbers
@@
The following lemma is useful for proving Theorem 1.2, which is essentially given in [33] (see equation (2.2) and (3.16) in [33]).
Lemma 2.1.
$(L_{\boldsymbol{A}})^{2}\,\,\leq\,\,\frac{1}{2}\sum_{j=1}^{m}\|\boldsymbol{a}_{j}\|_{2}^{2}-\frac{1}{2m\cdot\sin\frac{\pi}{2m}}\sum_{j=1}^{m}\|\boldsymbol{a}_{j}\|_{2}^{2}.$
Inspired by the proof of Theorem 3.3 in [33], we present the following lemma, which provides an alternative expression for the condition number of tight frames.
@@
$L_{\boldsymbol{B}}=\inf_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})\neq 0\end{subarray}}\frac{\||\boldsymbol{B}\boldsymbol{x}|-|\boldsymbol{B}\boldsymbol{y}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})}\\ \geq\frac{1}{\|\boldsymbol{M}^{\frac{1}{2}}\|_{2}}\cdot\inf_{\begin{subarray}{c}\widehat{\boldsymbol{x}},\widehat{\boldsymbol{y}}\in\mathbb{H}^{d}\\ \operatorname{dist}_{\mathbb{H}}(\widehat{\boldsymbol{x}},\widehat{\boldsymbol{y}})\neq 0\end{subarray}}\frac{\||\boldsymbol{A}\widehat{\boldsymbol{x}}|-|\boldsymbol{A}\widehat{\boldsymbol{y}}|\|_{2}}{\operatorname{dist}_{\mathbb{H}}(\widehat{\boldsymbol{x}},\widehat{\boldsymbol{y}})}=\frac{L_{\boldsymbol{A}}}{\|\boldsymbol{M}^{\frac{1}{2}}\|_{2}}.$
@@
We arrive at our conclusion. ∎
2.3. The perimeter-maximizing isodiametric problem
To prove our main stability results for phase retrieval, we introduce the perimeter-maximizing isodiametric problem, first studied by Reinhardt (1922) and later developed in a substantial literature (e.g., [26, 31, 14, 15, 16, 25, 9]). A convex polygon in $\mathbb{R}^{2}$ with $m$ edges is called a convex $m$-gon. The diameter of a polygon $P$, denoted by $\operatorname{diam}({P})$, is the largest distance between any pair of its vertices, i.e., $\operatorname{diam}({P}):=\max_{\boldsymbol{x},\boldsymbol{y}\,\in{P}}{\|\boldsymbol{x}-\boldsymbol{y}\|_{2}}$. The perimeter of a polygon $P$, denoted by $\operatorname{perim}({P})$, is the sum of the length of its edges. Denote r(P)r({P}) as the diameter-to-perimeter ratio of a polygon P{P}, i.e.,
$r(P):=\frac{\mathrm{diam}(P)}{\mathrm{perim}(P)}.$ (2.7)
The perimeter-maximizing isodiametric problem asks:
Problem 2.1.
Among all convex $m$-gons with fixed diameter, which ones have the maximal perimeter? Equivalently, which convex $m$-gons minimize the diameter-to-perimeter ratio $r(P)$?
@@ 2.3. The perimeter-maximizing isodiametric problem
  1. (i)
    Optimal polygon: A convex $m$-gon is defined as optimal if it is a solution to Problem 2.1. We say that the solution to Problem 2.1 is unique if it is unique up to scaling, translations, rotations, reflections, or any combination of these transformations.
  2. (ii)
  3. (iii)
  4. (iv)
  5. (v)
It is known that the optimal convex $m$-gon must be strictly convex [26]. Theorem 2.3 summarizes the existing results for the perimeter-maximizing isodiametric problem.
Theorem 2.3.
  1. (i)
    [26, 14, 16, 15, 31, 25] Let $m\geq 3$ be a positive integer that has an odd factor. Then
    $\min_{{P}\in\mathcal{P}_{m}}r({P})=\frac{1}{2m\cdot\sin\frac{\pi}{2m}},$
    where equality is achieved by finitely many equilateral strictly convex $m$-gons.
  2. (ii)
    [26, 14, 16, 15] Let $m=2^{s}$, where $s$ is an integer and $s\geq 2$. Then
    $\min_{{P}\in\mathcal{P}_{m}}r({P})>\frac{1}{2m\cdot\sin\frac{\pi}{2m}}.$
  3. (iii)
    [26, 14, 16, 15, 31] The regular $m$-gon $P_{m}$ is optimal if and only if $m$ is odd. Furthermore, ${P}_{m}$ is uniquely optimal if and only if $m$ is a prime.
In the following we introduce several basic properties of convex polygons, which will be utilized in later analysis.
@@
$\operatorname{diam}({P})=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}h({P},\boldsymbol{u})+h({P},-\boldsymbol{u}),$
where $h({P},\boldsymbol{z}):=\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{z}\rangle$ for any $\boldsymbol{z}\in\mathbb{S}^{1}$.
@@
$\operatorname{diam}(P)$ $=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}h(P,\boldsymbol{u})+h(P,-\boldsymbol{u})=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}(\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{u}\rangle+\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},-\boldsymbol{u}\rangle)$ (2.11)
$=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}(\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{u}\rangle-\min_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{u}\rangle)=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}\,{\langle\boldsymbol{v}_{j_{2}}-\boldsymbol{v}_{j_{1}},\boldsymbol{u}\rangle}.$
@@
$\langle\boldsymbol{v}_{j_{2}}-\boldsymbol{v}_{j_{1}},\boldsymbol{u}\rangle=\langle\sum_{j=j_{1}}^{j_{2}-1}{\boldsymbol{e}_{j}},\boldsymbol{u}\rangle\overset{(a)}{=}\langle\frac{1}{2}\sum_{j=j_{1}}^{j_{2}-1}{\boldsymbol{e}_{j}}-\frac{1}{2}\sum_{j=1}^{j_{1}-1}{\boldsymbol{e}_{j}}-\frac{1}{2}\sum_{j=j_{2}}^{m}{\boldsymbol{e}_{j}},\boldsymbol{u}\rangle\\ \overset{(b)}{=}\frac{1}{2}\sum_{j=1}^{m}{|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|},$ (2.13)
@@
$\langle\boldsymbol{e}_{j_{1}},\boldsymbol{u}\rangle=\langle\boldsymbol{v}_{j_{1}+1},\boldsymbol{u}\rangle-\langle\boldsymbol{v}_{j_{1}},\boldsymbol{u}\rangle\geq 0\quad\text{and}\quad\langle\boldsymbol{e}_{j_{2}-1},\boldsymbol{u}\rangle=\langle\boldsymbol{v}_{j_{2}},\boldsymbol{u}\rangle-\langle\boldsymbol{v}_{j_{2}-1},\boldsymbol{u}\rangle\geq 0.$ (2.14)
@@
$\langle\boldsymbol{e}_{t-1},\boldsymbol{u}\rangle<0,\quad\text{and}\quad\langle\boldsymbol{e}_{t},\boldsymbol{u}\rangle\geq 0,$ (2.15)
since otherwise we have $\langle\boldsymbol{e}_{i},\boldsymbol{u}\rangle<0$ for each $s+1\leq i\leq j_{2}-1$, which contradicts with (2.14). Since $P$ is strictly convex, for any $\boldsymbol{x}\in P$ there exist $\gamma_{1},\gamma_{2}\geq 0$ depending on $\boldsymbol{x}$ such that
@@
Hence, we arrive at our conclusion. This completes the proof. ∎
@@
[22] Suppose $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}\in\mathbb{R}^{2}$ are unit vectors that span $\mathbb{R}^{2}$, and suppose that $\alpha_{1},\dotsc,\alpha_{m}>0$. Then there exists a convex $m$-gon P{P} in $\mathbb{R}^{2}$, having edge unit outer normals $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}$ and corresponding edge lengths $\alpha_{1},\dotsc,\alpha_{m}>0$, if and only if $\sum_{j=1}^{m}\alpha_{j}\boldsymbol{u}_{j}=\boldsymbol{0}.$ Moreover, such a polygon P{P} is unique up to translation.
The next statement seems to be well known, but we could not find a precise reference. For completeness, we include a proof.
@@
Assume that $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$ are nonzero vectors with pairwise distinct directions and $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$. For all $j\in[m]$, rotating each vector $\frac{\boldsymbol{e}_{j}}{\|\boldsymbol{e}_{j}\|_{2}}$ counterclockwise by $\frac{\pi}{2}$ yields $\boldsymbol{u}_{j}$, while preserving the zero-sum property $\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}\boldsymbol{u}_{j}=\boldsymbol{0}$. It is easy to verify that $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}\in\mathbb{R}^{2}$ are unit vectors that span $\mathbb{R}^{2}$. Therefore we can utilize Theorem 2.4 to obtain a convex $m$-gon P{P} with edge unit outer normals $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}$ and corresponding edge lengths $\|\boldsymbol{e}_{1}\|_{2},\dotsc,\|\boldsymbol{e}_{m}\|_{2}$, i.e., with edge vectors $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$. Moreover, $P$ is unique, disregarding the translation of polygons. Now we prove that the convex polygon $P$ is strictly convex. Suppose, by contradiction, that $P$ is convex but not strictly convex. Then $P$ must have an interior angle of $\pi$, implying the existence of two edge vectors in the same direction. This contradicts the assumption that $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$ have pairwise distinct directions. Thus, $P$ is strictly convex.
Conversely, assume that P{P} is a strictly convex $m$-gon with the edge set $E$. Assume that $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$ are arranged counterclockwise. It is clear that $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$ and each $\boldsymbol{e}_{i}$ is nonzero. Without loss of generality, it is enough to prove that $\boldsymbol{e}_{1}$ and $\boldsymbol{e}_{i}$ have distinct directions for any $i>1$. For each $i\in\{2,\ldots,m\}$, define $\alpha_{i}$ as the counterclockwise angle from $\boldsymbol{e}_{1}$ to $\boldsymbol{e}_{i}$. Since $P$ is strictly convex, we have $\alpha_{i}\in(0,2\pi)$, and $\alpha_{i}$ is strictly increasing as $i$ increases. Thus, there does not exist an integer $i\in\{2,\ldots,m\}$ such that $\alpha_{i}=0$, i.e., $\boldsymbol{e}_{1}$ and $\boldsymbol{e}_{j}$ have the same direction. This completes the proof. ∎
3. Relationship between optimal tight frames and optimal polygons: Proof of Theorem 1.1
In this section, we aim to prove Theorem 1.1. We begin by introducing a bijection from $\mathcal{T}$ to $\mathbb{R}^{2}$.
@@
Let $f:\mathcal{T}\to\mathbb{R}^{2}$ be a map defined by
$f(\boldsymbol{a})\,\,=\,\,t^{2}\cdot(\cos 2\phi,\sin 2\phi)^{T}$ (3.1)
for any $\boldsymbol{a}=t\cdot(\cos\phi,\sin\phi)^{T}\in\mathcal{T}$, where $\mathcal{T}$ is defined in (1.7).
Remark 3.1.
The vector $f(\boldsymbol{a})$ is called the diagram vector associated with $\boldsymbol{a}$ [19]. We now show that the map $f:\mathcal{T}\to\mathbb{R}^{2}$ is a bijection. First, it holds that $\boldsymbol{a}=\boldsymbol{0}$ if and only if $f(\boldsymbol{a})=\boldsymbol{0}$. Second, for any nonzero vector $\boldsymbol{e}_{0}\in\mathbb{R}^{2}$, it possesses a unique representation in the form $\boldsymbol{e}_{0}=t_{0}\cdot(\cos{2\phi_{0}},\sin{2\phi_{0}})^{T}$, where $t_{0}>0$ and $\phi_{0}\in(0,\pi]$. Given this, the unique pre-image $\boldsymbol{a}_{0}\in\mathcal{T}$ such that $f(\boldsymbol{a}_{0})=\boldsymbol{e}_{0}$ is given by $\boldsymbol{a}_{0}=\sqrt{t_{0}}\cdot(\cos{\phi_{0}},\sin{\phi_{0}})^{T}$. Therefore, $f$ is both injective and surjective, implying it is a bijection. This ensures that the inverse map $f^{-1}:\mathbb{R}^{2}\to\mathcal{T}$ is well-defined, and is explicitly given by $f^{-1}(\boldsymbol{e}_{0})=\boldsymbol{a}_{0}$.
In the following theorem, we establish a bijective correspondence between optimal tight frames with $m$ vectors in $\mathbb{R}^{2}$ and optimal convex $m$-gons. This will play a key role in proving Theorem 1.1. We postpone the proof of Theorem 3.1 to the end of this section.
Theorem 3.1.
Assume that $\mathcal{A}=\{\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}\}\subset\mathcal{T}\subset{\mathbb{R}}^{2}$, where $\mathcal{T}$ is defined in (1.7). The set $\mathcal{A}$ is an optimal tight frame if and only if $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ is the edge vector set of an optimal $m$-gon $P\in\mathcal{P}_{m}$, where $f(\cdot)$ is defined in (3.1). In either case, we have
$\beta_{\boldsymbol{A}}=\frac{1}{\sqrt{1-2\cdot r({P})}},$
Remark 3.2.
Take $m=4$ as an example. Let $P\subset\mathbb{R}^{2}$ be the convex $4$-gon with the edge set $\{\boldsymbol{e}_{1},\boldsymbol{e}_{2},\boldsymbol{e}_{3},\boldsymbol{e}_{4}\}$, where
$\boldsymbol{e}_{1}=\begin{pmatrix}1\\ 0\end{pmatrix},\quad\boldsymbol{e}_{2}=\begin{pmatrix}\cos\frac{4\pi}{3}\\ \sin\frac{4\pi}{3}\end{pmatrix},\quad\boldsymbol{e}_{3}=2\sin\frac{\pi}{12}\begin{pmatrix}\cos\frac{3\pi}{4}\\ \sin\frac{3\pi}{4}\end{pmatrix},\quad\boldsymbol{e}_{4}=2\sin\frac{\pi}{12}\begin{pmatrix}\cos\frac{7\pi}{12}\\ \sin\frac{7\pi}{12}\end{pmatrix}$.
It is known that $P$ is the unique optimal 44-gon, and we have $r(P)=\frac{1}{2+\sqrt{6}-\sqrt{2}}$ [30, 29]. By Theorem 3.1 we see that the matrix $\boldsymbol{A}=[f^{-1}(\boldsymbol{e}_{1}),\ldots,f^{-1}(\boldsymbol{e}_{4})]^{T}\in\mathbb{R}^{4\times 2}$ has the minimal condition number $\beta_{\boldsymbol{A}}=\sqrt{1+\frac{\sqrt{2}}{2}+\frac{\sqrt{6}}{2}}\approx 1.71$. The optimal convex $4$-gon $P$ and the rows of $\boldsymbol{A}$ are both illustrated in Figure 1.
With the help of Lemma 2.3, Theorem 2.3 and Theorem 3.1, we present a proof of Theorem 1.1.
Proof of Theorem 1.1.
By Lemma 2.3 we have
3.1. Proof of Theorem 3.1
In order to establish Theorem 3.1, we first present Lemma 3.1 as a preliminary step. It has been proven in [19] that $\{\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}\}\subset{\mathbb{R}}^{2}$ forms a tight frame if and only if the diagram vectors $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ sum to zero, yielding a closed polygon in $\mathbb{R}^{2}$. Lemma 3.1 further clarifies the relationship between tight frames in $\mathbb{R}^{2}$ consisting of nonzero vectors with pairwise distinct directions and strictly convex polygons. Moreover, it connects the condition number of the matrix associated with such a frame to the diameter-to-perimeter ratio of the corresponding polygon.